OpenStudy (anonymous):

Okay, i am seriously confused with these questions. Solve each equation for x. 1. mx + nx = p 2. S = C + xC ( x/b is basicallly saying 1 over 2, like a fraction.) x + 2/y - 1 = 2

OpenStudy (anonymous):

1. x=p/(m+n) 2. x=(s/c)-1 3. x=2y-4